The average precipitation for the first 7 months of the year is 19.32 inches with a standard deviation of 2.4 inches. Assume that the average precipitation is normally distributed.

a. What is the probability that a randomly selected year will have precipitation greater than 18 inches for the first 7 months?

b. What is the average precipitation of 5 randomly selected years for the first 7 months?

c. What is the probability of 5 randomly selected years will have an average precipitation greater than 18 inches for the first 7 months?





1
Expert's answer
2022-03-09T04:00:56-0500

We are given "\mu" = 19.32 and "\sigma" = 2.4

a.

P(X>18) = P(X-"\mu" > 18-19.32) = P("\frac{X-\mu}{\sigma}" > "\frac{18 - 19.32}{2.4}")

We compute the z score using the formula:

z = "\frac{X-\mu}{\sigma}"

= "\frac{18-19.32}{2.4}"

= -0.55

P(X>18) = P(Z> -0.55)

where

P(Z> -0.55) = 1- P(Z<-0.55)

From the normal distribution table

P(Z<-0.55) = 0.2912

Therefore;

P(Z> -0.55) = 1- 0.2912

= 0.7088

Answer: 0.7088


b.


Answer:

We assume that the average precipitation of 5 randomly selected years for the first 7 months is the population mean "\mu" = 19.32


c.

Since "\mu" = 19.32 and "\sigma" = 2.4

We are given that the sample size n = 5

Now, we will use the formula:

z = "\frac{X-\mu}{\sigma/\sqrt{n}}"

P(X>18) = P(X-"\mu" > 18-19.32) = P("\frac{X-\mu}{\sigma/\sqrt{n}}" > "\frac{18 - 19.32}{2.4/\sqrt{5}}")

We compute the z score using the formula:

z = "\frac{X-\mu}{\sigma/\sqrt{n}}"

= "\frac{18-19.32}{2.4/\sqrt{5}}"

= -1.23

P(X>18) = P(Z> -1.23)

where

P(Z> -1.23) = 1- P(Z<-1.23)

From the normal distribution table

P(Z<-1.23) = 0.10935

Therefore;

P(Z> -1.23) = 1- 0.10935

= 0.89065


Answer: 0.8907